evaluate the sum A = 1 + 3 + 5 + ... + 2009 + 2011
Found 2 solutions by CubeyThePenguin, MathTherapy:
Answer by CubeyThePenguin(3113)

1 + 3 + ... + 2011 = (1006)^2 = 1,012,036

Answer by MathTherapy(10556)


evaluate the sum A = 1 + 3 + 5 + ... + 2009 + 2011
Yet ANOTHER WRONG answer by the other person!
Correct answer: